OpenStudy (anonymous):

What is the value of 4^−4? @hero Can you show me how to work this; please? :)

OpenStudy (anonymous):

@Hero

OpenStudy (solomonzelman):

\(\LARGE\color{red}{ \bf a^{-b}= }\)\(\huge\color{red}{ \bf \frac{1}{a^{b}} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 5^{-3}= }\)\(\huge\color{blue}{ \bf \frac{1}{5^3}= \frac{1}{125} }\)

OpenStudy (anonymous):

So in this case it would be 1/256?

OpenStudy (solomonzelman):

Yes, correct !

OpenStudy (anonymous):

Thanks! :)